# Nonclassical microwave radiation from the parametric dynamical Casimir effect in the reversed-dissipation regime of circuit optomechanics

### v1

@@ II Physical Model and The Effective System Hamiltonian
The Hamiltonian $\hat{H}_{m}$ is the free Hamiltonian of the MO and the three terms in $\hat{H}_{\rm pr}(\hat{H}_{\rm aux})$ denote, respectively, the free Hamiltonian of the primary (auxiliary) mode, the optomechanical coupling of the primary (auxiliary) mode to the MO with coupling strength $g_{0}=g_{0}^{\prime}/2$, and driving the primary (auxiliary) mode with the classical laser field of frequency $\omega_{L}(\omega_{L}^{\rm aux})$ and amplitude $E_{L}(E_{L}^{\rm aux})$. Here, an important point to remark is that the Hamiltonian of Eq. (3) consists of the extra term $g_{0}(\hat{a}^{\dagger 2}+\hat{a}^{2})(\hat{b}+\hat{b}^{\dagger})$, which is induced by the mechanical motion and responsible for the traditional DCE Macri2018 . However, in the most common experimental situation of cavity optomechanics, where the mechanical frequency is much smaller than the cavity frequency, this term can be neglected with good approximation. In fact, the experimental generation of a significant number of real photons from vacuum, sufficient to allow detection, requires the mechanical frequency to be at least twice that of the cavity, which still remains a serious problem. Recently a high mechanical frequency as large as $\omega_{m}/2\pi\sim 6$GHz has been reported Connell2010 ; Rouxinol2016 using microwave resonators and ultra-high-frequency mechanical micro- or nano-resonators. However, for the generation of Casimir radiation at the frequency of about  [^5] GHz a still higher mechanical frequency, $\omega_{m}/2\pi\sim 10$GHz is needed. As mentioned before to bypass this experimental difficulty, alternative schemes based on the simulation of the high-frequency mechanical motion can be exploited to produce the parametric DCE. This is the case addressed in the present paper, where the parametric DCE arises from the modulation of the driving laser frequency.
@@
$\hat{H}=-\Delta_{c}\hat{a}^{\dagger}\hat{a}+\omega_{m}\hat{b}^{\dagger}\hat{b}-g_{0}\hat{a}^{\dagger}\hat{a}(\hat{b}+\hat{b}^{\dagger})+iE_{L}(\hat{a}^{\dagger}-\hat{a}),$ (5)
@@
$\hat{F}_{\rm in}(t)=\hat{a}_{\rm in}+\sqrt{\frac{\Gamma_{m}}{\kappa}}\frac{g_{0}\hat{a}}{\omega_{m}}(\hat{\tilde{b}}_{\rm in}^{\dagger}-\hat{\tilde{b}}_{\rm in}),$ (10)
@@
$+\frac{4}{\kappa_{\rm aux}}+\absolutevalue{\frac{\kappa_{\rm aux}/2-2i\Delta_{\rm aux}}{\kappa^{2}_{\rm aux}/4+4\Delta_{\rm aux}^{2}}}^{2}+\frac{2\sqrt{\kappa_{\rm aux}}}{\kappa^{2}_{\rm aux}/4+4\Delta_{\rm aux}^{2}}\Big{)}\Big{]},$ (11)
@@ II.1 Simulation of the parametric DCE Hamiltonian
By using $[\hat{a},\hat{a}^{\dagger}]=1$ and $[\hat{a},\sqrt{\hat{n}}]=\frac{1}{2\sqrt{\hat{n}}}\hat{a}$, it is straightforward to show that the deformed operators $\hat{A}$ and $\hat{A}^{\dagger}$ obey the (deformed) commutation relation
$[\hat{A},\hat{A}^{\dagger}]=1+C_{K}(2n+\frac{1}{4}e^{1/n}),$ (19)
@@
As discussed in literature Law1994 ; Dodonov2011 ; Dodonov1996 ; Saito2002 ; Nation2012 a single resonant cavity mode with an externally prescribed time-dependent frequency can be used as a paradigm for understanding the mechanism underlying the photon generation. Actually, harmonic time-dependence of the cavity eigenfrequency is analogous to an apparent periodic displacement of the cavity mirrors which is responsible for parametric amplification of the cavity field mode in the DCE. Within the framework of instantaneous mode functions and the associated dynamical Fock space Law1994 , the dynamics of the cavity field in the absence of dissipation can be effectively described by the Hamiltonian
@@
$+iC_{\epsilon}(t)C_{K}(\hat{a}^{\dagger}\sqrt{\hat{n}}\hat{a}^{\dagger}\sqrt{\hat{n}}-\sqrt{\hat{n}}\hat{a}\sqrt{\hat{n}}\hat{a})\Big{]}$
@@
$+iC_{\epsilon}(t)C_{K}(\hat{a}^{\dagger}\sqrt{\hat{n}}\hat{a}^{\dagger}\sqrt{\hat{n}}-\sqrt{\hat{n}}\hat{a}\sqrt{\hat{n}}\hat{a})\big{]}.$
@@
$+iC_{\epsilon}(t)C_{K}(\hat{a}^{\dagger}\sqrt{\hat{n}}\hat{a}^{\dagger}\sqrt{\hat{n}}-\sqrt{\hat{n}}\hat{a}\sqrt{\hat{n}}\hat{a})\Big{]}.$
@@ II.2 Numerical solution
$\dot{\hat{\rho}}=$ $i[\hat{\rho},\hat{H}_{\rm DCE(WCR)}]+\frac{\kappa}{2}(\hat{a}\hat{\rho}\hat{a}^{\dagger}-\hat{a}^{\dagger}\hat{a}\hat{\rho}-\hat{\rho}\hat{a}^{\dagger}\hat{a}),$ (28)
where the Hamiltonians $\hat{H}_{\rm DCE(WCR)}$ are given in Eqs. (26), and (27), respectively. By using the Quantum Toolbox in Python (QuTip) qutip1 ; qutip2 , we numerically solve the master equation to obtain the time-evolved density matrix of the system $\hat{\rho}(t)$ by which we can evaluate the mean number of the generated microwave Casimir photons and their quantum statistical properties. Note that solving the master equation with Hamiltonian (26) enables us to obtain the exact dynamics of the system beyond the WCR.
@@ II.3 Analytical solution
×$\times{\rm exp}\left[\frac{\alpha(t)}{2}\hat{a}^{{\dagger}2}\right]{\rm exp}\left[\frac{\beta(t)}{2}\hat{n}\right]{\rm exp}\left[\frac{\gamma(t)}{2}\hat{a}^{2}\right],$
@@ III Casimir Photon Generation
Figure 2: Time evolution of the mean number of generated Casimir photons $n_{\rm Casimir}=\bra{0}\hat{n}(t)\ket{0}$ based on the exact numerical solution of the master equation (28) with Hamiltonian $\hat{H}_{\rm DCE}$ in Eq. (26) for different values of the scaled Kerr nonlinearity parameter $C_{K}$. The other system parameters are set as $\omega_{m}/2\pi=5.33\text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $\widetilde{C}_{\epsilon}=10^{-1}$, $C_{E}=10^{-2}$, $\kappa/2\pi=118\text{kHz}$.
Figure 3: Time evolution of the mean number of generated Casimir photons $n_{\rm Casimir}(t)=\langle 0|\hat{n}(t)|0\rangle$ in the WCR, obtained from both the analytical expression in Eq. (31) and numerical simulations based on Hamiltonian (27). The orange dashed line shows the evolution in the absence of dissipation. The system parameters are set as $\omega_{m}/2\pi=5.33~\text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $C_{E}=10^{-2}$, and $\kappa/2\pi=118~\text{kHz}$. Here, the Kerr nonlinearity strength is fixed to $C_{K}=10^{-3}$ while $\widetilde{C}_{\epsilon}$ is varied to explore its influence on photon generation.
@@
In order to see more clearly the role played in the Casimir photon creation by the time modulation of the driving laser frequency and the induced Kerr nonlinearity, we consider the rate of Casimir photon production. From the perspective of practical applications, notably in free-space or in fiber, the photon production rate is an important quantity which can be experimentally measured by using single-photon avalanche detectors (SPADs) vezzoli2019 . In the system under investigation, the average rates of the Casimir photon production corresponding to Figs. 2(a), 2(b), 3(a), and 3(b) are, respectively, obtained as 10Mcps, 50Mcps, 40Mcps, and 1.23Gcps (count or photon per second) . Accordingly, we understand that when the induced Kerr nonlinearity is extremely weak, the rate of Casimir photon creation increases significantly as the modulation amplitude parameter increases a little bit. The high generation rate in range from MHz to GHz implies that the proposed scheme can be seen as a high-brightness nonclassical microwave quantum source. (The nonclassicality of the generated Casimir radiation will be verified in the following Secs. IVV, and VI.)
@@ IV Photon Counting Statistics
Figure 5: Time evolution of the Mandel parameter in the WCR, obtained from both the analytical expression in Eq. (34) and numerical simulations based on Hamiltonian (27). The orange dashed line shows the evolution in the absence of dissipation. The system parameters are set as $\omega_{m}/2\pi=5.33~\text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $C_{E}=10^{-2}$, and $\kappa/2\pi=118~\text{kHz}$, while $\widetilde{C}_{\epsilon}$ is varied to explore its influence on photon counting statistics.
We now focus on the quantum statistical properties of the generated Casimir photons, particularly their photon counting statistics as described by the Mandel parameter. This quantity indicates whether the photon statistics is sub-Poissonian, Poissonian, or super-Poissonian, and thus provides useful insight into the quantum or classical nature of the emitted radiation. The Mandel parameter is defined as mandel1986
@@
As can be seen from Fig. 4(a), for $C_{K}<\widetilde{C}_{\epsilon}$ , the Mandel parameter decreases towards negative values and reaches $Q=-0.5$ in the early stages of the system evolution, indicating the nonclassical sub-Poissonian statistics of the generated Casimir photons. Then, after making damped oscillations, it is finally stabilized at an asymptotical zero value corresponding to the Poissonian statistics. On the other hand, Figs. 4(b)-4(c) reveal that as $C_{K}$ increases towards values exceeding $\widetilde{C}_{\epsilon}$, the Mandel parameter remains positive at all times which corresponds to the classical super-Poissonian statistics.
@@
In view of the generation mechanism of quadrature squeezed states in one hand, and the presence of the term proportional to $\hat{a}^{2\dagger}-\hat{a}^{2}$ in the Hamiltonians of Eqs. (26) and (27) on the other hand, it is reasonable to expect that the generated Casimir photons exhibit the squeezing effect. To explore the squeezing property of the generated Casimir photons, we define the dimensionless quadrature amplitudes $\hat{q}=\frac{\hat{a}+\hat{a}^{\dagger}}{2}$ and $\hat{p}=\frac{\hat{a}-\hat{a}^{\dagger}}{2i}$ related to the position and momentum operators of the cavity mode, respectively. A quantum state of the cavity mode is said to be squeezed when one of the quadrature components $\hat{q}$ and $\hat{p}$ satisfies the relation $\expectationvalue{(\delta\hat{O})^{2}}<1/4(\hat{O}=\hat{q}\text{ or }\hat{p})$. The degree of quadrature squeezing can be quantified in dB (decibel) unit via
@@
[…]
@@
Now, we examine the temporal behaviour of $S_{O}(t)$ which gives information on the squeezing of the quadrature $O(O=q,p)$. In Figs. 6(a)-(d) we have plotted the squeezing parameters $S_{q}(t)$ (left panels) and $S_{p}(t)$ (right panels) for the generated Casimir photons with respect to time, obtained from the numerical solution of the master equation (28) with the Hamiltonian [Eq. (26)]. For panels (a) and (b) we take the Kerr coefficient $C_{K}=0.05$ and for panels (c) and (d) we choose $C_{K}=0.1$. The modulation amplitude parameter $\widetilde{C}_{\epsilon}$ is fixed at value 0.10.10.1. It is seen that both $S_{q}(t)$ and $S_{p}(t)$ oscillate as functions of time, showing alternatively some degree of quadrature squeezing (up to 4.3 dB) at short times, but any squeezing disappears as time passes due to the cavity field dissipation. We also find that with increasing the Kerr coefficient, the quadratures squeezing can reappear at longer times, albeit with smaller amplitudes. Figs. 6(e)-(h) demonstrate the time evolution of $S_{q}(t)$ (left panels) and $S_{p}(t)$ (right panels) calculated by using the analytical expressions in Eqs. (36) (green-solid curves) as well as the numerical simulation of the master equation (28) with the Hamiltonian $\hat{H}_{\rm WCR}$ in Eq. (27), with and without considering the cavity dissipation (red-solid and orange-dashed curves, respectively). As can be seen, in the WCR and with the same cavity dissipation rate as used in the numerical analysis with the Hamiltonian $\hat{H}_{\rm DCE}$ (red-solid curves), a higher quadrature squeezing (up to 5dB) can be achieved. More importantly, the quadratures squeezing can persist over a relatively longer time, indicating an improved resilience to dissipation. These results suggest that the WCR described by Eq. (27) provides better performance and robustness against environmental losses compared to the full effective Hamiltonian $\hat{H}_{\rm DCE}$ in Eq. (26).

### v2

@@ II Physical Model and The Effective System Hamiltonian
The Hamiltonian $\hat{H}_{m}$ is the free Hamiltonian of the MO and the three terms in $\hat{H}_{\rm pr}(\hat{H}_{\rm aux})$ denote, respectively, the free Hamiltonian of the primary (auxiliary) mode, the optomechanical coupling of the primary (auxiliary) mode to the MO with coupling strength $g_{0}=g_{0}^{\prime}/2$, and driving the primary (auxiliary) mode with the classical laser field of frequency $\omega_{L}(\omega_{L}^{\rm aux})$ and amplitude $E_{L}(E_{L}^{\rm aux})$. Here, an important point to remark is that the Hamiltonian of Eq. (3) consists of the extra term $g_{0}(\hat{a}^{\dagger 2}+\hat{a}^{2})(\hat{b}+\hat{b}^{\dagger})$, which is induced by the mechanical motion and responsible for the traditional DCE [101]. However, in the most common experimental situation of cavity optomechanics, where the mechanical frequency is much smaller than the cavity frequency, this term can be neglected with good approximation. In fact, the experimental generation of a significant number of real photons from vacuum, sufficient to allow detection, requires the mechanical frequency to be at least twice that of the cavity, which still remains a serious problem. Recently a high mechanical frequency as large as $\omega_{m}/2\pi\sim 6$GHz has been reported [60, 126] using microwave resonators and ultra-high-frequency mechanical micro- or nano-resonators. However, for the generation of Casimir radiation at the frequency of about 55GHz a still higher mechanical frequency, $\omega_{m}/2\pi\sim 10$GHz is needed. As mentioned before to bypass this experimental difficulty, alternative schemes based on the simulation of the high-frequency mechanical motion can be exploited to produce the parametric DCE. This is the case addressed in the present paper, where the parametric DCE arises from the modulation of the driving laser frequency.
@@
[…]
@@
$\hat{F}_{\rm in}(t)=\hat{a}_{\rm in}-\sqrt{\frac{\Gamma_{m}}{\kappa}}\frac{g_{0}\hat{a}}{\omega_{m}}(\hat{\tilde{b}}_{\rm in}^{\dagger}-\hat{\tilde{b}}_{\rm in}),$ (10)
@@
[…]
@@
It is important to note that the Kerr-type term in Eq.(12) is fundamentally different from the standard Kerr interaction in conventional $\chi^{(3)}$ media. In ordinary Kerr materials, the nonlinearity is fixed by intrinsic material properties and is therefore not tunable. In our case, the Kerr interaction arises effectively from the adiabatic elimination of the mechanical mode in the RDR of optomechanics, making its strength $g_{K}$ fully controllable through optomechanical coupling, which constitutes an advantage not available in standard Kerr platforms.
@@ II.1 Simulation of the parametric DCE Hamiltonian
By using $\hat{a},\hat{a}^{\dagger}]=1$ and $\hat{a},\sqrt{\hat{n}}]=\frac{1}{2\sqrt{\hat{n}}}\hat{a}$, it is straightforward to show that the deformed operators $\hat{A}$ and $\hat{A}^{\dagger}$ obey the (deformed) commutation relation
$\hat{A},\hat{A}^{\dagger}]=1+C_{K}(2n+\frac{1}{4}e^{1/n}),$ (19)
@@
As discussed in literature [127, 13, 142, 129, 5] a single resonant cavity mode with an externally prescribed time-dependent frequency can be used as a paradigm for understanding the mechanism underlying the photon generation. Actually, harmonic time-dependence of the cavity eigenfrequency is analogous to an apparent periodic displacement of the cavity mirrors which is responsible for parametric amplification of the cavity field mode in the DCE. Within the framework of instantaneous mode functions and the associated dynamical Fock space [127], the dynamics of the cavity field in the absence of dissipation can be effectively described by the Hamiltonian
@@
[…]
@@
[…]
@@
[…]
@@ II.2 Numerical solution
[…]
where the Hamiltonians $\hat{H}_{\rm DCE(WCR)}$ are given in Eqs. (26), and (27), respectively. By using the Quantum Toolbox in Python (QuTip) [130, 131], we numerically solve the master equation to obtain the time-evolved density matrix of the system $\hat{\rho}(t)$ by which we can evaluate the mean number of the generated microwave Casimir photons and their quantum statistical properties. Note that solving the master equation with Hamiltonian (26) enables us to obtain the exact dynamics of the system beyond the WCR.
It is important to mention that our numerical analysis must rely on the effective Hamiltonians in Eqs. (26) and (27), as a direct simulation of the optomechanical Hamiltonian in Eq. (5) is not practically feasible. In the instantaneous-mode quantization required for time-dependent cavity boundaries, the operators and the effective detuning acquire explicit time dependence, making the operator basis itself time-dependent and unsuitable for standard numerical methods. Applying the canonical transformation of Law’s dynamical-Casimir formalism [127] removes this difficulty by mapping the instantaneous operators to a fixed mode and transferring all time dependence to the Hamiltonian parameters. This yields the effective Hamiltonian in Eq. (21), from which Eqs. (26) and (27) follow and which provide the only numerically tractable formulation for our simulations.
@@ II.3 Analytical solution
$\times{\rm exp}\left[\frac{\alpha(t)}{2}\hat{a}^{{\dagger}2}\right]{\rm exp}\left[\frac{\beta(t)}{2}\hat{n}\right]{\rm exp}\left[\frac{\gamma(t)}{2}\hat{a}^{2}\right],$
@@ III Casimir Photon Generation
[…]
Figure 3: Time evolution of the mean number of generated Casimir photons $n_{\rm Casimir}(t)=\langle 0|\hat{n}(t)|0\rangle$ in the WCR, obtained from both the analytical expression in Eq. (31) and numerical simulations based on Hamiltonian (27). The orange dashed line shows the evolution in the absence of dissipation. The system parameters are set as $\omega_{m}/2\pi=5.33\penalty 10000\ \text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $C_{E}=10^{-2}$, and $\kappa/2\pi=118\penalty 10000\ \text{kHz}$. Here, the Kerr nonlinearity strength is fixed to $C_{K}=10^{-3}$ while $\widetilde{C}_{\epsilon}$ is varied to explore its influence on photon generation.
@@
In order to see more clearly the role played in the Casimir photon creation by the time modulation of the driving laser frequency and the induced Kerr nonlinearity, we consider the rate of Casimir photon production. From the perspective of practical applications, notably in free-space or in fiber, the photon production rate is an important quantity which can be experimentally measured by using single-photon avalanche detectors (SPADs) [41]. In the system under investigation, the average rates of the Casimir photon production corresponding to Figs. 2(a), 2(b), 3(a), and 3(b) are, respectively, obtained as 10Mcps, 50Mcps, 40Mcps, and 1.23Gcps (count or photon per second) . Accordingly, we understand that when the induced Kerr nonlinearity is extremely weak, the rate of Casimir photon creation increases significantly as the modulation amplitude parameter increases a little bit. The high generation rate in range from MHz to GHz implies that the proposed scheme can be seen as a high-brightness nonclassical microwave quantum source. (The nonclassicality of the generated Casimir radiation will be verified in the following Secs. IV, V, and VI.)
@@ IV Photon Counting Statistics
Figure 5: Time evolution of the Mandel parameter in the WCR, obtained from both the analytical expression in Eq. (34) and numerical simulations based on Hamiltonian (27). The orange dashed line shows the evolution in the absence of dissipation. The system parameters are set as $\omega_{m}/2\pi=5.33\penalty 10000\ \text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $C_{E}=10^{-2}$, and $\kappa/2\pi=118\penalty 10000\ \text{kHz}$, while $\widetilde{C}_{\epsilon}$ is varied to explore its influence on photon counting statistics.
We now focus on the quantum statistical properties of the generated Casimir photons, particularly their photon counting statistics as described by the Mandel parameter. This quantity indicates whether the photon statistics is sub-Poissonian, Poissonian, or super-Poissonian, and thus provides useful insight into the quantum or classical nature of the emitted radiation. The Mandel parameter is defined as [132]
@@
As can be seen from Fig. 4(a), for $C_{K}<\widetilde{C}_{\epsilon}$ , the Mandel parameter decreases towards negative values and reaches Q=0.5Q=-0.5 in the early stages of the system evolution, indicating the nonclassical sub-Poissonian statistics of the generated Casimir photons. Then, after making damped oscillations, it is finally stabilized at an asymptotical zero value corresponding to the Poissonian statistics. On the other hand, Figs. 4(b)-4(c) reveal that as $C_{K}$ increases towards values exceeding $\widetilde{C}_{\epsilon}$, the Mandel parameter remains positive at all times which corresponds to the classical super-Poissonian statistics.
@@
[…]
@@
with $\expectationvalue{(\delta\hat{O})^{2}}_{\rm vac}=1/4$ as the quantum-vacuum fluctuation. Then, the condition for squeezing in the quadrature component can be simply written as $S_{O}>0$. Since 3-dB squeezing, which corresponds to $50\%$ noise reduction below the zero-point level (i.e., $\expectationvalue{(\delta\hat{O})^{2}}=\expectationvalue{(\delta\hat{O})^{2}}_{\rm vac}/2=1/8$) is a limit for many proposals, the squeezing beyond 3 dB can be regarded as strong squeezing. In the WCR, by using the Hamiltonian (26), the variances of the cavity mode quadratures fluctuations can be obtained as (see Appendix D).
@@
Now, we examine the temporal behaviour of $S_{O}(t)$ which gives information on the squeezing of the quadrature $O(O=q,p)$. In Figs. 6(a)-(d) we have plotted the squeezing parameters $S_{q}(t)$ (left panels) and $S_{p}(t)$ (right panels) for the generated Casimir photons with respect to time, obtained from the numerical solution of the master equation (28) with the Hamiltonian [Eq. (26)]. For panels (a) and (b) we take the Kerr coefficient $C_{K}=0.05$ and for panels (c) and (d) we choose $C_{K}=0.1$. The modulation amplitude parameter $\widetilde{C}_{\epsilon}$ is fixed at value $0.1$. It is seen that both $S_{q}(t)$ and $S_{p}(t)$ oscillate as functions of time, showing alternatively some degree of quadrature squeezing (up to 4.3 dB) at short times, but any squeezing disappears as time passes due to the cavity field dissipation. We also find that with increasing the Kerr coefficient, the quadratures squeezing can reappear at longer times, albeit with smaller amplitudes. Figs. 6(e)-(h) demonstrate the time evolution of $S_{q}(t)$ (left panels) and $S_{p}(t)$ (right panels) calculated by using the analytical expressions in Eqs. (36) (green-solid curves) as well as the numerical simulation of the master equation (28) with the Hamiltonian $\hat{H}_{\rm WCR}$ in Eq. (27), with and without considering the cavity dissipation (red-solid and orange-dashed curves, respectively). As can be seen, in the WCR and with the same cavity dissipation rate as used in the numerical analysis with the Hamiltonian $\hat{H}_{\rm DCE}$ (red-solid curves), a higher quadrature squeezing (up to 5dB) can be achieved. More importantly, the quadratures squeezing can persist over a relatively longer time, indicating an improved resilience to dissipation. These results suggest that the WCR described by Eq. (27) provides better performance and robustness against environmental losses compared to the full effective Hamiltonian $\hat{H}_{\rm DCE}$ in Eq. (26).